The following information pertains to Marvolo, Inc. What is the sales volume (in units) required to obtain a target AFTER-TAX profit of $108,000?

Selling price per unit

$100

Variable costs per unit

$75

Total fixed costs

$425,000

Tax rate

40%
Answer #1

Contribution margin=Sales-Variable cost  

=100-75=$25 per unit

Target before tax profits=108,000/(1-tax rate)

=108,000/(1-0.4)=$180,000

Target Contribution margin=Target profits+Fixed costs

=(425,000+180,000)=$605,000

Hence target sales=605,000/25

=24200 units